2x2−4y2−20x−16y+26=0
|
Find the Standard Form of the Hyperbola
Answer


Answer by stanbon(75887)

2x2−4y2−20x−16y+26=0
|
Find the Standard Form of the Hyperbola
---------------------
2x^2 - 20x -4y^2 - 16y = -26
-----
2(x^2-10x) - 4(y^2+4y) = -26
-----
Complete the square::
2(x^2-10x+25) -4(y^2+4y+4) = -26 + 2*25 - 4*4
----
2(x-5)^2 -4(y+2)^2 = 8
-----
(x-5)^2/4 - (y+2)^2/2 = 1
================================
Cheers,
Stan H.
----------------